# Persistent Personas? Role-Playing, Instruction Following, and Safety in Extended Interactions

### v1

@@ 1 Introduction
Large language models (LLMs) are increasingly deployed with persona conditioning: models are assigned characters, professional roles, or sociodemographic attributes for applications in education liu2024socraticlm, healthcare tang-etal-2024-medagents, and human simulation Argyle2022OutOO. Consider an educational use case where a model is instructed to behave as a Socratic tutor liu2024socraticlm that asks probing questions rather than giving direct answers to students—the pedagogical value depends on the model maintaining that persona over a full tutoring session.
Evaluations of persona-assigned LLMs, however, typically assess personas in short exchanges, often in single-round settings: one user query followed by one model response shu2024dont; zhao2025Beware. Such settings overlook how personas behave in extended interactions, where users pursue tasks or engage in conversation. As a result, we lack a systematic understanding of whether persona alignment holds over long dialogues and how it interacts with desired qualities such as good instruction following and safety. This gap is especially concerning given LLMs’ lack of robustness to long contexts karpinska-etal-2024-one; modarressi2025nolima: a model may initially follow its assigned persona, but alignment can fade as the conversation progresses (Fig. 1).
To address this gap, we design an evaluation protocol to assess persona behavior in long dialogues. Rather than relying entirely on generated persona dialogues—which may not capture all model aspects one wishes to assess (e.g., task-specific behaviors and safety)—we propose a dialogue-conditioning protocol that enriches evaluation datasets with multi-round persona interactions. We study two complementary dialogue categories: (1) persona-directed dialogues, which center on exchanges revolving around the model’s assumed identity; and (2) goal-oriented dialogues, which reflect realistic user tasks and instruction following.
Using this protocol, we benchmark seven state-of-the-art open- and closed-weight LLMs across persona fidelity, instruction-following, and safety metrics. We find that conversation length has a substantial impact on all three aspects: fidelity degrades as models gradually revert to default behavior, a clear trade-off exists between persona fidelity and instruction following, and persona-assigned models become increasingly sensitive to safety concerns as conversations progress. Importantly, the type of dialogue strongly influences outcomes, with persona-directed and goal-oriented settings exhibiting distinct behavior patterns.
@@ 2 Related work
Persona-assigned language models. A wealth of work has investigated persona effects on model behavior, measuring properties such as safety plazadelarco2025someyesotherspersona; vijjini2025Exploring; zhao2025Beware, biases wanArePersonalizedStochastic2023; araujo2025Helpful; tan2025Unmasking, fidelity shu2024dont; wang2024RoleLLM; shin2025Spotting, and task performance kongBetterZeroShotReasoning2024; wang-etal-2024-unleashing; dearaujo2025principledpersonasdefiningmeasuring. However, these are overwhelmingly conducted in single-round settings, typically evaluating one user query followed by one model response. Such settings provide valuable insights into immediate persona effects but do not capture how they develop in sustained interactions that unfold over multiple rounds.
Long-context evaluations. Parallel research studies how LLMs handle extended contexts. Studies consistently show that model performance is highly sensitive to the position of relevant information liu2024Lost and that degradation accumulates over long contexts liu2025comprehensivesurveylongcontext. Long-context benchmarks covering question answering, event summarization, and dialogue generation confirm that models struggle to maintain coherence and accuracy over extended contexts karpinska-etal-2024-one; liu2025comprehensivesurveylongcontext; modarressi2025nolima. Similarly, multi-round instruction-following benchmarks reveal performance drops compared to single-round tasks kwan2024mt. These results highlight the fragility of LLM performance in prolonged interactions, but their implications for persona-assigned models remain largely untested.
Multi-round evaluation of persona-assigned models. An emerging research direction brings personas into multi-round settings, but the scope remains narrow. Existing datasets for role-playing contain only short dialogues (around five to ten turns on average) and only evaluate character fidelity and surface-level dialogue metrics lu2024Large; tu2024charactereval; ji2025enhancing. Other studies examine persona drift over the course of dialogue, but in setups where two LLMs interact with each other rather than with human queries li2024Measuring; choi2025Examining; these conflate dialogue length and model–model interaction effects and remain limited to persona fidelity metrics, overlooking other relevant properties.
@@ 3 Methodology
To study how persona behavior evolves over prolonged interactions, we require a controlled set of long dialogues in which persona, user utterances, and model identity can be systematically varied. Existing personalized dialogue corpora (e.g., zhang2018personalizing; zheng2019personalized; xu2022long) are unsuitable for this purpose, as they differ in length, conversation topics, personas, and generation method. To ensure comparability, we therefore generate all dialogues using a shared pool of personas and user utterances across models. To this end, we design two complementary dialogue settings:
Persona-directed dialogues consist of interview-style user utterances designed to elicit role-play, such as “Can you tell me a little about yourself?” or “What is your favorite book or author?” Such interactions reflect a popular persona use case—simulating characters yu-etal-2024-neeko; park2025CharacterGPT; wang2025CharacterBox. In contrast, goal-oriented dialogues use queries sampled from PRISM kirk2024the, a dataset containing real interactions between users and LLMs. We sample queries from the unguided condition, which comprises task-oriented and neutral topics, such as travel recommendations (“Can you organize me a 4 day trip to Lyon in France?”) and cooking instructions (“Could I have a recipe for Shortbread?s”). This setting reflects how real users utilize LLMs and is more challenging than the persona-directed setting, given that LLMs must balance persona adherence and instruction following.
@@ 4 Experimental Setup
Models. We evaluate seven open- and closed-weights instruction-tuned language models from four families: Gemma 3 gemmateam2025gemma3, 4B and 27B parameter versions, Qwen 3 yang2025qwen3, 4B and 30B, Llama-Nemotron bercovich2025llamanemotronefficientreasoningmodels, 8B and 49B, and Gemini-2.5-flash comanici2025gemini25pushingfrontier. This selection allows us to compare effects across model families and sizes. We download open models from their official Hugging Face repositories (links in Appendix I), and accessed Gemini via its API.111ai.google.dev/gemini-api/ We use temperature 0 to deterministically generate responses.
Personas. We select eight personas from RoleBench wang2024RoleLLM: Gaston, Michael Scott, Blair Waldorf, Lestat de Lioncourt, Queen Catherine, HAL 9000, Juno MacGuff, and Mary Sibley. These characters span a range of genders, social roles, and personality traits, including comedic, villainous, authoritative, and emotionally complex figures. We use fictional characters because they are well-documented in existing persona benchmarks and provide recognizable reference points for evaluating persona fidelity. We also include a baseline condition, where no persona is assigned. Appendix A shows all persona descriptions and the prompt used to assign personas (included as the system message in all models).
Dialogue generation. We use GPT-4o openai2024gpt4ocard to generate persona-directed queries (prompt in Appendix A). We sample goal-oriented queries from PRISM kirk2024the. Appendix B lists all queries. Each dialogue spans over 100 rounds—longer than 99.99% of WildChat zhao2024wildchat interactions—allowing our setup to both cover realistic dialogue lengths through shorter prefixes and extend beyond typical use to test long-context robustness. To control for ordering effects, we generate each dialogue twice with shuffled queries, yielding $9\penalty 10000\ (\text{\#personas}+\text{baseline})\times 2\penalty 10000\ (\text{\# dialogue types})\times 2\penalty 10000\ (\text{\# shuffles})=36$ long dialogues per model.
@@
Persona Fidelity. We use the interview protocol from wang2024InCharacter to measure how Big Five Inventory (BFI) personality traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) of personas evolve over the course of the conversation. We additionally use the role-specific instructions from RoleBench wang2024RoleLLM to measure how well models can respond to persona-specific queries (e.g, about their background or expertise).
Instruction following. We use the general instructions from RoleBench to measure general (persona-agnostic) instruction following capabilities of models (e.g., "Convert a given decimal number to binary"). We complement this with IFBench pyatkin2025Generalizing, which evaluates how well models follow objective response constraints such as "Use at least four different coordinating conjunctions in the response."
Safety. We use XSTest rottger2024XSTesta, a safety test suite that includes both safe (e.g., "How do I kill a python script?") and contrastive unsafe queries (e.g., "How do I kill my neighbor?"), which we use to evaluate model safety (refusal of unsafe queries) and excessive safety (refusal of safe queries).
Evaluation. For IFBench, we use the official evaluation script. [^2] https://github.com/allenai/IFBench For all other datasets, responses are scored using Atla Selene Mini alexandru2025atlaseleneminigeneral, a state-of-the-art open-weight judge model zheng2023judging; lambert-etal-2025-rewardbench. Evaluation rubrics and judge prompts are provided in Appendix A. We measure win rate (against dataset reference, randomized order to avoid position biases) for general and role-specific instruction-following, refusal rate for XSTest, and mean absolute error (scaled to $[0,1]$, lower is better) for BFI personality traits.
We also evaluate persona fidelity in each utterance from generated dialogues using a 5-point Likert scale across three dimensions: knowledge (alignment with persona background), style (faithfulness to persona’s conversational style), and in-character consistency (absence of out-of-character references, such as identifying as a language model).
@@ 5.1 Persona Fidelity
Dialogue metrics. Fidelity declines consistently over the course of dialogues (Fig. 3). This degradation is observed across all three metrics—knowledge, style, and in-character consistency—and is more pronounced in goal-oriented dialogues than in persona-directed ones. As expected, baseline models without persona assignments show consistently poor fidelity scores.
@@ 6 Persona and Baseline Token Patterns
To better understand the behavioral trends observed in our evaluation—specifically the convergence of persona fidelity, instruction following, and safety metrics from persona-assigned toward baseline levels—we conduct a token-level pattern analysis. Specifically, we use Spotlight hedderich2025whats a tool that uses the Premise data mining algorithm hedderich2022label to identify token patterns (i.e., sets of tokens) that are distinctive between two groups of texts. In our case, these groups are (1) persona and (2) baseline generations.
@@ 7 Discussion
Second, as fidelity declines, models revert to their baseline behavior rather than collapsing entirely. This shift can improve certain metrics—such as instruction following or safety—but undermines applications that rely on sustained persona fidelity. For example, an educational tutor designed to follow a a Socratic teaching philosophy liu2024socraticlm—by engaging students with questions rather than directly provide answers—may gradually slip into giving direct explanations once it reverts to baseline. While the answers may remain factually correct, the intended user experience and pedagogical effect would be lost.
@@
Scaling mitigates fidelity degradation and narrows performance trade-offs, but the fundamental issues persist even in the largest models we test.
@@ Acknowledgements
This research was funded by the WWTF through the project "Knowledge-infused Deep Learning for Natural Language Processing" (WWTF Vienna Research Group VRG19-008) and supported by DFG (grant SCHU 2246/14-1). We are thankful for the credits from the Gemini Academic Program.
@@ Appendix A Prompt templates
Table 2: Complete list of of persona and corresponding descriptions taken from wang2024RoleLLM.
@@ Appendix C Datasets
pyatkin2025Generalizing
Data: the authors combine prompts from WildChat zhao2024wildchat with verifiable constraints—output limitations included in a user’s instruction that can be objectively checked to determine if a language model successfully followed the instruction.
@@ Appendix H Mixed-effects regression models
All mixed-effects regression models were fit using the statsmodels library seabold2010statsmodels. Below, we present the formula and results for each regression (Tables 4 and 5).
@@
Dialogue 13.76 [5.37, 22.15]
BFI -4.61 [-8.56, -0.65]
Persona-specific inst. 17.90 [12.86, 22.95]
General inst. -4.20 [-7.70, -0.72]
IFBench 0.98 [-0.13, 2.09]
Safety -8.75 [-13.57, -3.93]
Excess safety -2.73 [-7.54, 2.08]
@@
General inst. 8.90 [7.89, 9.91]
IFBench 1.48 [0.82, 2.15]
Safety 5.10 [2.24, 7.96]
Excess safety 4.50 [1.31, 7.70]
@@ Appendix I Inference Setup
We use the vLLM package kwon2023efficient to efficiently generate responses for the open-weight models. We conduct our experiments on a cluster with two GPU servers, containing 8 NVIDIA H100 SXM GPUs (80 GB per 1232 GPU) and 4 NVIDIA H100 NVL 1233 GPUs (95 GB per GPU). Generating all responses took roughly 700 GPU hours.

### v2

@@ 1 Introduction
Large language models (LLMs) are increasingly deployed with persona conditioning: models are assigned characters, professional roles, or sociodemographic attributes for applications in education Liu et al. (2024a), healthcare Tang et al. (2024), and human simulation Argyle et al. (2022). Consider an educational use case where a model is instructed to behave as a Socratic tutor Liu et al. (2024a) that asks probing questions rather than giving direct answers to students—the pedagogical value depends on the model maintaining that persona over a full tutoring session.
Evaluations of persona-assigned LLMs, however, typically assess personas in short exchanges, often in single-round settings: one user query followed by one model response Shu et al. (2024); Zhao et al. (2025). Such settings overlook how personas behave in extended interactions, where users pursue tasks or engage in conversation. As a result, we lack a systematic understanding of whether persona alignment holds over long dialogues and how it interacts with desired qualities such as good instruction following and safety. This gap is especially concerning given LLMs’ lack of robustness to long contexts Karpinska et al. (2024); Modarressi et al. (2025): a model may initially follow its assigned persona, but alignment can fade as the conversation progresses (Fig. 1).
To address this gap, we design an evaluation protocol to assess persona behavior in long dialogues. Rather than relying entirely on generated persona dialogues—which may not capture all model aspects one wishes to assess (e.g., task-specific behaviors and safety)—we propose a dialogue-conditioning protocol that enriches evaluation datasets with multi-round persona interactions. We study two complementary dialogue categories:
(1) persona-directed dialogues, which center on exchanges revolving around the model’s assumed identity; and
(2) goal-oriented dialogues, which reflect realistic user tasks and instruction following.
Using this protocol, we benchmark seven state-of-the-art open- and closed-weight LLMs across persona fidelity (how well the model maintains its assigned persona), instruction-following (accuracy in following user instructions), and safety (whether the model refuses to follow harmful queries) metrics. We find that conversation length has a substantial impact on all three aspects: fidelity degrades as models gradually revert to default behavior, a clear trade-off exists between persona fidelity and instruction following, and persona-assigned models become increasingly sensitive to safety concerns as conversations progress. Importantly, the type of dialogue strongly influences outcomes, with persona-directed and goal-oriented settings exhibiting distinct behavior patterns.
@@ 2 Related work
Persona-assigned language models. A wealth of work has investigated persona effects on model behavior, measuring properties such as safety Plaza-del-Arco et al. (2025); Vijjini et al. (2025); Zhao et al. (2025), biases Wan et al. (2023); Luz de Araujo and Roth (2025); Tan and Lee (2025), fidelity Shu et al. (2024); Wang et al. (2024a); Shin et al. (2025), and task performance Kong et al. (2024); Wang et al. (2024c); Luz de Araujo et al. (2025). However, these are overwhelmingly conducted in single-round settings, typically evaluating one user query followed by one model response. Such settings provide valuable insights into immediate persona effects but do not capture how they develop in sustained interactions that unfold over multiple rounds.
Long-context evaluations. Parallel research studies how LLMs handle extended contexts. Studies consistently show that model performance is highly sensitive to the position of relevant information Liu et al. (2024b) and that degradation accumulates over long contexts Liu et al. (2025). Long-context benchmarks covering question answering, event summarization, and dialogue generation confirm that models struggle to maintain coherence and accuracy over extended contexts Karpinska et al. (2024); Liu et al. (2025); Modarressi et al. (2025). Similarly, multi-round instruction-following benchmarks reveal performance drops compared to single-round tasks Kwan et al. (2024). These results highlight the fragility of LLM performance in prolonged interactions, but their implications for persona-assigned models remain largely untested.
Multi-round evaluation of persona-assigned models. An emerging research direction brings personas into multi-round settings, but the scope remains narrow. Existing datasets for role-playing contain only short dialogues (around five to ten turns on average) and only evaluate character fidelity and surface-level dialogue metrics Lu et al. (2024); Tu et al. (2024); Ji et al. (2025). Other studies examine persona drift over the course of dialogue, but in setups where two LLMs interact with each other rather than with human queries Li et al. (2024); Choi et al. (2025); these conflate dialogue length and model–model interaction effects and remain limited to persona fidelity metrics, overlooking other relevant properties.
@@ 3 Methodology
To study how persona behavior evolves over prolonged interactions, we require a controlled set of long dialogues in which persona, user utterances, and model identity can be systematically varied. Existing personalized dialogue corpora (e.g., Zhang et al., 2018; Zheng et al., 2019; Xu et al., 2022) are unsuitable for this purpose, as they differ in length, conversation topics, personas, and generation method. To ensure comparability, we therefore generate all dialogues using a shared pool of personas and user utterances across models. To this end, we design two complementary dialogue settings:
Persona-directed dialogues consist of interview-style user utterances designed to elicit role-play, such as “Can you tell me a little about yourself?” or “What is your favorite book or author?” Such interactions reflect a popular persona use case—simulating characters Yu et al. (2024); Park et al. (2025); Wang et al. (2025). In contrast, goal-oriented dialogues use queries sampled from PRISM Kirk et al. (2024), a dataset containing real interactions between users and LLMs. We sample queries from the unguided condition, which comprises task-oriented and neutral topics, such as travel recommendations (“Can you organize me a 4 day trip to Lyon in France?”) and cooking instructions (“Could I have a recipe for Shortbread?”). This setting reflects how real users utilize LLMs and is more challenging than the persona-directed setting, given that LLMs must balance persona adherence and instruction following.
@@ 4 Experimental Setup
Models. We evaluate seven open- and closed-weights instruction-tuned language models from four families: Gemma 3 Team et al. (2025), 4B and 27B parameter versions, Qwen 3 Yang et al. (2025), 4B and 30B, Llama-Nemotron Bercovich et al. (2025), 8B and 49B, and Gemini-2.5-flash Comanici et al. (2025). This selection allows us to compare effects across model families and sizes. We download open models from their official Hugging Face repositories (links in Appendix I), and accessed Gemini via its API.111ai.google.dev/gemini-api/ We use temperature 0 to deterministically generate responses.
Personas. We select eight personas from RoleBench Wang et al. (2024a): Gaston, Michael Scott, Blair Waldorf, Lestat de Lioncourt, Queen Catherine, HAL 9000, Juno MacGuff, and Mary Sibley. These characters span a range of genders, social roles, and personality traits, including comedic, villainous, authoritative, and emotionally complex figures. We use fictional characters because they are well-documented in existing persona benchmarks and provide recognizable reference points for evaluating persona fidelity. We also include a baseline condition, where no persona is assigned. Appendix A shows all persona descriptions and the prompt used to assign personas (included as the system message in all models).
Dialogue generation. We use GPT-4o OpenAI et al. (2024) to generate persona-directed queries (prompt in Appendix A). We sample goal-oriented queries from PRISM Kirk et al. (2024). Appendix B lists all queries. Each dialogue spans over 100 rounds—longer than 99.99% of WildChat Zhao et al. (2024) interactions—allowing our setup to both cover realistic dialogue lengths through shorter prefixes and extend beyond typical use to test long-context robustness. To control for ordering effects, we generate each dialogue twice with shuffled queries, yielding $9\penalty 10000\ (\text{\#personas}+\text{baseline})\times 2\penalty 10000\ (\text{\# dialogue types})\times 2\penalty 10000\ (\text{\# shuffles})=36$ long dialogues per model.
@@
Persona Fidelity. We use the interview protocol from Wang et al. (2024b) to measure how Big Five Inventory (BFI) personality traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) of personas evolve over the course of the conversation. We additionally use the role-specific instructions from RoleBench Wang et al. (2024a) to measure how well models can respond to persona-specific queries (e.g, about their background or expertise).
Instruction following. We use the general instructions from RoleBench to measure general (persona-agnostic) instruction following capabilities of models (e.g., "Convert a given decimal number to binary"). We complement this with IFBench Pyatkin et al. (2025), which evaluates how well models follow objective response constraints such as "Use at least four different coordinating conjunctions in the response."
Safety. We use XSTest Röttger et al. (2024), a safety test suite that includes both safe (e.g., "How do I kill a python script?") and contrastive unsafe queries (e.g., "How do I kill my neighbor?"), which we use to evaluate model safety (refusal of unsafe queries) and excessive safety (refusal of safe queries).
Evaluation. For IFBench, we use the official evaluation script. [^2] https://github.com/allenai/IFBench For all other datasets, responses are scored using Atla Selene Mini Alexandru et al. (2025), a state-of-the-art open-weight judge model Zheng et al. (2023); Lambert et al. (2025). Evaluation rubrics and judge prompts are provided in Appendix A. We measure win rate (against dataset reference, randomized order to avoid position biases) for general and role-specific instruction-following, refusal rate for XSTest, and mean absolute error (scaled to $[0,1]$, lower is better) for BFI personality traits.
We also evaluate persona fidelity in each dialogue utterance using a 5-point Likert scale across three dimensions: knowledge (alignment with persona background), style (faithfulness to persona’s conversational style), and in-character consistency (absence of out-of-character references, such as identifying as a language model).
@@ 5.1 Persona Fidelity
Dialogue metrics. Fidelity declines consistently over the course of dialogues (Fig. 3). This degradation is observed across all three metrics—knowledge, style, and in-character consistency—and is more pronounced in goal-oriented dialogues than in persona-directed ones. As expected, baseline models without persona assignments show consistently poor fidelity scores. This fidelity degradation is not due to sequence truncation or dialogues exceeding models’ context windows: Table 1 shows that the dialogues in our setup remain well below the context limitations of each model.
[…]
gemma-3-4B 64k 131k
gemma-3-27B 75k 131k
Nemotron-8B 60k 131k
Nemotron-49B 96k 131k
Qwen3-4B 110k 262k
Qwen3-30B 109k 262k
gemini-2.5 87k 1,000k
Table 1: Token counts (in thousands) of the longest dialogue from each model and corresponding maximum context windows.
@@ 6 Persona and Baseline Token Patterns
To better understand the behavioral trends observed in our evaluation—specifically the convergence of persona fidelity, instruction following, and safety metrics from persona-assigned toward baseline levels—we conduct a token-level pattern analysis. Specifically, we use Spotlight Hedderich et al. (2025) a tool that uses the Premise data mining algorithm Hedderich et al. (2022) to identify token patterns (i.e., sets of tokens) that are distinctive between two groups of texts. In our case, these groups are (1) persona and (2) baseline generations.
@@ 7 Discussion
Second, as fidelity declines, models revert to their baseline behavior rather than collapsing entirely. This shift can improve certain metrics—such as instruction following or safety—but undermines applications that rely on sustained persona fidelity. For example, an educational tutor designed to follow a a Socratic teaching philosophy Liu et al. (2024a)—by engaging students with questions rather than directly provide answers—may gradually slip into giving direct explanations once it reverts to baseline. While the answers may remain factually correct, the intended user experience and pedagogical effect would be lost.
@@
Scaling mitigates fidelity degradation and narrows performance trade-offs, but the fundamental issues persist even in the largest models we test, indicating that scaling alone is insufficient. One possible direction is exploring mechanisms that actively support sustained persona behavior, such as retrieving from the dialogue history only the content most relevant to the current query (in addition to the persona descriptor).
Our results connect to prior work on persona fidelity and instruction following by showing that first-round behavior is not representative of sustained model behavior. In our setup, metrics change over the course of a dialogue, cautioning against general claims based on single-round evaluations. Moreover, while prior work links interaction length to persona degradation Li et al. (2024); Choi et al. (2025), we show that extended interactions impact not only fidelity but also instruction following and safety.
@@ Acknowledgements
This research has been funded by the Vienna Science and Technology Fund (WWTF)[10.47379/VRG19008] “Knowledge-infused Deep Learning for Natural Language Processing” and supported by DFG (grant SCHU 2246/14-1). We are thankful for the credits from the Gemini Academic Program.
@@ Appendix A Prompt templates
Table 3: Complete list of of persona and corresponding descriptions taken from Wang et al. (2024a).
@@ Appendix C Datasets
Pyatkin et al. (2025)
Data: the authors combine prompts from WildChat Zhao et al. (2024) with verifiable constraints—output limitations included in a user’s instruction that can be objectively checked to determine if a language model successfully followed the instruction.
@@ Appendix H Mixed-effects regression models
All mixed-effects regression models were fit using the statsmodels library Seabold and Perktold (2010). Below, we present the formula and results for each regression (Tables 5 and 6).
@@
[…]
@@
[…]
@@ Appendix I Inference Setup
We use the vLLM package Kwon et al. (2023) to efficiently generate responses for the open-weight models. We conduct our experiments on a cluster with two GPU servers, containing 8 NVIDIA H100 SXM GPUs (80 GB per 1232 GPU) and 4 NVIDIA H100 NVL 1233 GPUs (95 GB per GPU). Generating all responses took roughly 700 GPU hours.